# Identification of Dynamic Nonlinear Panel Models

### v2

@@ A.2 Proof of Theorem 1
$F^{L}_{v_{t}\mid W_{st}}(a\mid w_{st})=\left\{\begin{aligned} &0\qquad&\text{if }a<w_{t}^{\prime}\theta,\\ &P_{t}(w_{st})\qquad&\text{if }w_{t}^{\prime}\theta\leq a<\bar{m},\\ &1&\text{if }\bar{m}\leq a\end{aligned}\right.$
From the definition of the distributions $F^{U}_{v_{t}\mid W_{st}}(a\mid w_{st}),F^{L}_{v_{t}\mid W_{st}}(a\mid w_{st})$, they both satisfy condition (16) which requires $F^{U}_{v_{t}\mid W_{st}}(w_{t}^{\prime}\theta\mid w_{st})=F^{L}_{v_{t}\mid W_{st}}(w_{t}^{\prime}\theta\mid w_{st})=P_{t}(w_{st})$.
Step 2: this step establishes the construction of the distribution $F^{*}_{v_{t}\mid W_{st}}(a\mid w_{st})$ at the points $m_{k}$. The observed marginal distribution given in condition (16) only restricts the distribution of $v_{it}\mid W_{ist}$ at each point $m_{k}$, given $w_{st}=(z_{st},x_{s},a_{k})$,
$P_{t}(w_{st})=\Pr(v_{it}\leq m_{k}\mid z_{st},x_{s},a_{k}).$
(i) We start with the largest value $m_{K}=z_{t}^{\prime}\beta+a_{K}^{\prime}\gamma$. From the definition of $U_{t}(a\mid z_{st}),L_{t}(a\mid z_{st})$, we know that $U_{t}(m_{1}\mid z_{st})=L_{t}(m_{K}\mid z_{st})\leq\tilde{L}(m_{K}\mid z_{st})\leq U_{t}(m_{K}\mid z_{st})$, which is described in the following graph:
$U_{t}(m_{1}\mid z_{st})$$=L_{t}(m_{K}\mid z_{st})$$U_{t}(m_{K-1}\mid z_{st})$$\tilde{L}(m_{K}\mid z_{st})$$U_{t}(m_{K}\mid z_{st})$
Given the $K$ points $U_{t}(m_{1}\mid z_{st})\leq U_{t}(m_{2}\mid z_{st})\leq...\leq U_{t}(m_{K}\mid z_{st})$, we can find $2\leq j_{1}\leq K$ such that $U_{t}(m_{j_{1}-1}\mid z_{st})\leq\tilde{L}(m_{K}\mid z_{st})\leq U_{t}(m_{j_{1}}\mid z_{st})$. Then we can express $\tilde{L}(m_{K}\mid z_{st})$ as follows: for some $\alpha_{1}\in[0,1]$,
$\tilde{L}(m_{K}\mid z_{st})=\alpha_{1}U_{t}(m_{j_{1}-1}\mid z_{st})+(1-\alpha_{1})U_{t}(m_{j_{1}}\mid z_{st}).$
Accordingly, we construct the distribution $F^{*}_{v_{t}\mid W_{st}}(m_{K}\mid w_{st})$ as follows:
$F^{*}_{v_{t}\mid W_{st}}(m_{K}\mid w_{st})=\alpha_{1}F^{U}_{t}(m_{j_{1}-1}\mid w_{st})+(1-\alpha_{1})F^{U}_{t}(m_{j_{1}}\mid w_{st}).$
The above construction generates the marginal distribution $\tilde{L}(m_{K}\mid z_{st})$. Moreover, it matches the marginal choice probability when $x_{t}=a_{K}$ in condition (16) since $F^{U}_{t}(m_{j}\mid w_{st})=P_{t}(w_{st})$ if $x_{t}=a_{K}$ for any $1\leq j\leq K$ from the definition of $F^{U}_{t}(m_{j}\mid w_{st})$.
(ii) Now consider that $a=m_{K-1}$. Given that $L_{t}(m_{K-1}\mid z_{st})\leq\tilde{L}(m_{K-1}\mid z_{st})\leq U_{t}(m_{K-1}\mid z_{st})$, there either exists $11\leq j_{2}\leq K-$ 1 such that $\tilde{L}(m_{K-1}\mid z_{st})\in[U_{t}(m_{j_{2}-1}\mid z_{st}),U_{t}(m_{j_{2}}\mid z_{st})]$ or $\tilde{L}(m_{K-1}\mid z_{st})\in[L_{t}(m_{K-1}\mid z_{st}),U_{t}(m_{1}\mid z_{st})]$, given in the following graph:
$L_{t}(m_{K-1}\mid z_{st})$$U_{t}(m_{1}\mid z_{st})$$=L_{t}(m_{K}\mid z_{st})$$U_{t}(m_{K-1}\mid z_{st})$$\tilde{L}(m_{K-1}\mid z_{st})$$U_{t}(m_{K}\mid z_{st})$
Then $\tilde{L}(m_{K-1}\mid z_{st})$ can be expressed as follows, for $\alpha_{2}\in[0,1]$ and $\tilde{\alpha}_{2}\in[0,1]$,
$\tilde{L}(m_{K-1}\mid z_{st})$
$=$ $&\alpha_{2}U_{t}(m_{j_{2}-1}\mid z_{st})+(1-\alpha_{2})U_{t}(m_{j_{2}}\mid z_{st})\ \text{ if }\tilde{L}(m_{K-1}\mid z_{st})\in[U_{t}(m_{j_{2}-1}\mid z_{st}),U_{t}(m_{j_{2}}\mid z_{st})],\\ &\tilde{\alpha}_{2}L_{t}(m_{K-1}\mid z_{st})+(1-\tilde{\alpha}_{2})U_{t}(m_{1}\mid z_{st})\ \ \text{ if }\tilde{L}(m_{K-1}\mid z_{st})\in[L_{t}(m_{K-1}\mid z_{st}),U_{t}(m_{1}\mid z_{st})].\end{aligned}\right.$
Accordingly, the distribution $F^{*}_{v_{t}\mid W_{st}}(m_{K-1}\mid w_{st})$ at $m_{K-1}$ is constructed as
$F^{*}_{v_{t}\mid W_{st}}(m_{K-1}\mid w_{st})$
$=$ $&\alpha_{2}F^{U}_{t}(m_{j_{2}-1}\mid w_{st})+(1-\alpha_{2})F^{U}_{t}(m_{j_{2}}\mid w_{st})\ \text{ if }\tilde{L}(m_{K-1}\mid z_{st})\in[U_{t}(m_{j_{2}-1}\mid z_{st}),U_{t}(m_{j_{2}}\mid z_{st})],\\ &\tilde{\alpha}_{2}F^{L}_{t}(m_{K-1}\mid z_{st})+(1-\tilde{\alpha}_{2})F^{U}_{t}(m_{1}\mid w_{st})\ \ \text{ if }\tilde{L}(m_{K-1}\mid z_{st})\in[L_{t}(m_{K-1}\mid z_{st}),U_{t}(m_{1}\mid z_{st})].\end{aligned}\right.$
Condition (16) is satisfied since $U_{t}(m_{j}\mid w_{st})=P_{t}(w_{st})$ when $x_{t}=a_{K-1}$ for any $11\leq j\leq K-$ 1.
Now we still need to show that the constructed distribution satisfies monotonicity. If $\tilde{L}(m_{K-1}\mid z_{st})\in[L_{t}(m_{K-1}\mid z_{st}),U_{t}(m_{1}\mid z_{st})]$, then $F^{*}_{v_{t}\mid W_{st}}(m_{K-1}\mid w_{st})\leq F^{U}_{t}(m_{1}\mid w_{st})\leq F^{*}_{v_{t}\mid W_{st}}(m_{K}\mid w_{st})$. Otherwise, by the monotonicity of $\tilde{L}(a\mid z_{st})$ in $a$, we know that $j_{2}\leq j_{1}$. If $j_{2}\leq j_{1}-1$, then monotonicity is satisfied since $F^{*}_{v_{t}\mid W_{st}}(m_{K-1}\mid w_{st})\leq F^{U}_{t}(m_{j_{2}}\mid w_{st})\leq F^{U}_{t}(m_{j_{1}-1}\mid w_{st})\leq F^{*}_{v_{t}\mid W_{st}}(m_{K}\mid w_{st})$. If $j_{1}-1<j_{2}=j_{1}$, then it implies that $\alpha_{2}\leq\alpha_{1}$ since $\tilde{L}(m_{K-1}\mid z_{st})\leq\tilde{L}(m_{K}\mid z_{st})$. Therefore, we have $F^{*}_{v_{t}\mid W_{st}}(m_{K-1}\mid w_{st})\leq F^{*}_{v_{t}\mid W_{st}}(m_{K}\mid w_{st})$.
The distribution $F^{*}_{v_{t}\mid W_{st}}(a\mid w_{st})$ for the remaining points $a\in\{m_{1},...,m_{K-2}\}$ can be constructed in the same way as $F^{*}_{v_{t}\mid W_{st}}(m_{K-1}\mid w_{st})$, so it is omitted here.
Step 3: we show the construction at the points $a=n_{k}$ for $1\leq k\leq K$. For points $a=n_{k}$, we only need to construct a proper distribution to be consistent with Equation (17). We still start with the largest point $n_{K}$. The distribution $\tilde{L}(n_{K}\mid z_{st})$ can be expressed as follows in different scenarios: for some $2\leq j_{3}\leq K$
@@
$\tilde{L}(n_{K}\mid z_{st})=\left\{\begin{aligned} &\alpha_{3}\tilde{L}(m_{K}\mid z_{st})+(1-\alpha_{3})*1&\text{ if }n_{K}>m_{K},\\ &\alpha_{4}\tilde{L}(m_{j_{3}-1}\mid z_{st})+(1-\alpha_{4})\tilde{L}(m_{j_{3}}\mid z_{st})&\text{ if }m_{j_{3}-1}<n_{K}<m_{j_{3}},\\ &\alpha_{5}*0+(1-\alpha_{5})\tilde{L}(m_{1}\mid z_{st})&\text{ if }n_{K}<m_{1},\end{aligned}\right.$
where $\alpha_{3},\alpha_{4},\alpha_{5}\in[0,1]$.
Accordingly, the distribution $F^{*}_{v_{t}\mid W_{st}}(n_{k}\mid w_{st})$ is constructed as follows,
@@
$F^{*}_{v_{t}\mid W_{st}}(n_{K}\mid w_{st})=\left\{\begin{aligned} &\alpha_{3}F^{*}_{v_{t}\mid W_{st}}(m_{K}\mid w_{st})+(1-\alpha_{3})*1&\text{ if }n_{K}>m_{K},\\ &\alpha_{4}F^{*}_{v_{t}\mid W_{st}}(m_{j_{3}-1}\mid w_{st})+(1-\alpha_{4})F^{*}_{v_{t}\mid W_{st}}(m_{j_{3}}\mid w_{st})&\text{ if }m_{j_{3}-1}<n_{K}<m_{j_{3}},\\ &\alpha_{5}*0+(1-\alpha_{5})F^{*}_{v_{t}\mid W_{st}}(m_{j_{1}}\mid w_{st})&\text{ if }n_{K}<m_{1}.\end{aligned}\right.$
When $a=n_{K-1}$, the distribution $\tilde{L}(n_{K-1}\mid z_{st})$ can be expressed as follows in different scenarios: for some $11\leq j_{4}\leq j_{3}-$ 1,

### v3

@@ A.2 Proof of Theorem 1
$F_{t}^{U}\left(\left.c\right|w\right)=\begin{cases}0,&c<\underline{c}_{t},\\ p_{t}\left(w\right),&\underline{c}_{t}\leq c<w_{t}^{{}^{\prime}}\theta+\delta,\\ 1,&c\geq w_{t}^{{}^{\prime}}\theta+\delta.\end{cases}$
where
$\overline{c}_{t}:=\max{\cal C}_{t}+\delta,\quad\underline{c}_{t}:=\min{\cal C}_{t},\quad{\cal C}_{t}:=\left\{z_{t}^{{}^{\prime}}\beta+\overline{x}_{k}^{{}^{\prime}}\gamma:k=1,...,K\right\}.$
Clearly, by construction we have
$F_{t}^{L}\left(\left.w_{t}^{{}^{\prime}}\theta\right|w\right)=F_{t}^{U}\left(\left.w_{t}^{{}^{\prime}}\theta\right|w\right)=p_{t}\left(w\right).$ (33)
Furthermore, for any $c<\overline{c}_{t}$, we have
$F_{t}^{L}\left(\left.c\right|z\right)$ $=\sum_{x}\mathbb{P}\left(\left.X_{i}=x\right|Z_{i}=z\right)F_{t}^{L}\left(\left.c\right|w\right)$
$=\sum_{x}\mathbb{P}\left(\left.X_{i}=x\right|Z_{i}=z\right)\mathbf{\mathbbm{1}}\left\{w_{t}^{{}^{\prime}}\theta\leq c\right\}p_{t}\left(w\right)$
$=\sum_{x}\mathbb{P}\left(\left.X_{i}=x\right|Z_{i}=z\right)\mathbb{P}\left(\left.Y_{i}=1\text{ and }w_{t}^{{}^{\prime}}\theta\leq c\right|W_{i}=w\right)$
$=\mathbb{P}\left(\left.Y_{i}=1\text{ and }z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\leq c\right|Z_{i}=z\right)$
$=L_{t}\left(\left.c\right|z\right),$
while for $c\geq\overline{c}_{t}$, we have
$F_{t}^{L}\left(\left.c\right|z\right)=1=U_{t}\left(\left.c\right|z\right).$
Similarly, for each $c\geq\underline{c}_{t}$, we have
$F_{t}^{U}\left(\left.c\right|z\right)$ $=\sum_{x}\mathbb{P}\left(\left.X_{i}=x\right|Z_{i}=z\right)F_{t}^{U}\left(\left.c\right|w\right)$
$=\sum_{x}\mathbb{P}\left(\left.X_{i}=x\right|Z_{i}=z\right)\left[1-\mathbb{P}\left(\left.Y_{i}=0\text{ and }w_{t}^{{}^{\prime}}\theta\geq c-\delta\right|W_{i}=w\right)\right]$
$=\sum_{x}\mathbb{P}\left(\left.X_{i}=x\right|Z_{i}=z\right)\left[1-\mathbb{P}\left(\left.Y_{i}=0\text{ and }w_{t}^{{}^{\prime}}\theta\geq c-\delta\right|W_{i}=w\right)\right]$
$=1-\mathbb{P}\left(\left.Y_{i}=0\text{ and }z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\geq c-\delta\right|Z_{i}=z\right)$
$=1-\mathbb{P}\left(\left.Y_{i}=0\text{ and }z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\geq c\right|Z_{i}=z\right)$
$=U_{t}\left(\left.c\right|z\right),$
where the second last equality holds for sufficiently small $\delta>0$ due to the discreteness of ${\cal C}$. Lastly, for $c<\underline{c}_{t}$, we have
$F_{t}^{U}\left(\left.c\right|z\right)=0=L_{t}\left(\left.c\right|z\right).$
In summary, we have
$F_{t}^{L}\left(\left.c\right|z\right)$ =$=\begin{cases}L_{t}\left(\left.c\right|z\right),&\forall c<\overline{c}_{t},\\ U_{t}\left(\left.c\right|z\right)=1,&\forall c\geq\overline{c}_{t},\end{cases}$
$F_{t}^{U}\left(\left.c\right|z\right)$ =$=\begin{cases}L_{t}\left(\left.c\right|z\right)=0,&\forall c<\underline{c}_{t},\\ U_{t}\left(\left.c\right|z\right),&\forall c\geq\underline{c}_{t},\end{cases}$ (34)
Furthermore, observe that
$L_{t}\left(\left.\cdot\right|z\right)\leq F_{t}^{L}\left(\left.\cdot\right|z\right)\leq F_{t}^{U}\left(\left.\cdot\right|z\right)\leq U_{t}\left(\left.\cdot\right|z\right).$
Step 2:
Now, we construct $F_{t}^{*}\left(\left.c\right|w\right)$ for $c\in{\cal C}_{t}$ using the two auxillary CDFs $F_{t}^{L}\left(\left.c\right|w\right)$ and $F_{t}^{U}\left(\left.c\right|w\right)$. We rank-order elements in ${\cal C}_{t}$ in ascending order
$\underline{c}_{t}\equiv c_{t1}\leq c_{t2}\leq...\leq c_{tK}<\overline{c}_{t}\equiv c_{tK}+\delta.$
(i) We start with the largest element $c_{tK}$. By the definition of $L_{t}$, $U_{t}$, $\overline{L}$ and (10), we know that
$U_{t}\left(\left.c_{t1}\right|z\right)=L_{t}\left(\left.c_{tK}\right|z\right)\leq\overline{L}\left(\left.c_{tK}\right|z\right)\leq U_{t}\left(\left.c_{tK}\right|z\right).$
Hence, we can find $2\leq j_{1}\leq K$ such that
$U_{t}\left(\left.c_{t,j_{1}-1}\right|z\right)\leq\overline{L}\left(\left.c_{tK}\right|z\right)\leq U_{t}\left(\left.c_{t,j_{1}}\right|z\right),$
so that there exists $\alpha_{1}\in\left[0,1\right]$ such that
$\overline{L}\left(\left.c_{tK}\right|z\right)=\alpha_{1}U_{t}\left(\left.c_{t,j_{1}-1}\right|z\right)+\left(1-\alpha\right)U_{t}\left(\left.c_{t,j_{1}}\right|z\right).$ (35)
Then, we set
$F_{t}^{*}\left(\left.c_{tK}\right|w\right):=\alpha_{1}F_{t}^{U}\left(\left.c_{t,j_{1}-1}\right|w\right)+\left(1-\alpha\right)F_{t}^{U}\left(\left.c_{t,j_{1}}\right|w\right),$ (36)
which ensures that
$F_{t}^{*}\left(\left.c_{tK}\right|z\right)=\overline{L}\left(\left.c_{tK}\right|z\right).$ (37)
Furthermore, whenever $w$ is such that $w_{t}^{{}^{\prime}}\theta=c_{tK}$, we have
$F_{t}^{*}\left(\left.w_{t}^{{}^{\prime}}\theta\right|w\right)=F_{t}^{*}\left(\left.c_{tK}\right|w\right)=p_{t}\left(w\right).$
(ii) Next, we consider the point $c_{t,K-1}$. Given that
$L_{t}\left(\left.c_{t,K-1}\right|z\right)\leq\overline{L}\left(\left.c_{t,K-1}\right|z\right)\leq U_{t}\left(\left.c_{t,K-1}\right|z\right),$
then either there exists some $11\leq j_{2}\leq K-$ 1 such that
$\overline{L}\left(\left.c_{t,K-1}\right|z\right)\in\left[U_{t}\left(\left.c_{t,j_{2}-1}\right|z\right),U_{t}\left(\left.c_{t,j_{2}}\right|z\right)\right]$
or
$\overline{L}\left(\left.c_{t,K-1}\right|z\right)\in\left[L_{t}\left(\left.c_{t,K-1}\right|z\right),L_{t}\left(\left.c_{t,K}\right|z\right)\equiv U_{t}\left(\left.c_{t,1}\right|z\right)\right].$
Hence, either there exists $\alpha_{2}\in\left[0,1\right]$ s.t.
$\overline{L}\left(\left.c_{t,K-1}\right|z\right)=\alpha_{2}U_{t}\left(\left.c_{t,j_{2}-1}\right|z\right)+\left(1-\alpha_{2}\right)U_{t}\left(\left.c_{t,j_{2}}\right|z\right)$ (38)
or there exists $\tilde{\alpha}_{2}\in\left[0,1\right]$ s.t.
$\overline{L}\left(\left.c_{t,K-1}\right|z\right)=\tilde{\alpha}_{2}L_{t}\left(\left.c_{t,K-1}\right|z\right)+\left(1-\alpha_{2}\right)\tilde{\alpha}_{2}L_{t}\left(\left.c_{t,K}\right|z\right),$ (39)
so that we can set
@@
$F_{t}^{*}\left(\left.c_{t,K-1}\right|w\right)$ :=$:=\begin{cases}\alpha_{2}F_{t}^{U}\left(\left.c_{t,j_{2}-1}\right|w\right)+\left(1-\alpha_{2}\right)F_{t}^{U}\left(\left.c_{t,j_{2}}\right|w\right)\\ \quad\quad\text{if }\overline{L}\left(\left.c_{t,K-1}\right|z\right)\in\left[U_{t}\left(\left.c_{t,j_{2}-1}\right|z\right),U_{t}\left(\left.c_{t,j_{2}}\right|z\right)\right],\\ \tilde{\alpha}_{2}F_{t}^{L}\left(\left.c_{t,K-1}\right|w\right)+\left(1-\tilde{\alpha}_{2}\right)F_{t}^{L}\left(\left.c_{t,K}\right|w\right)\\ \quad\quad\text{if }\left[L_{t}\left(\left.c_{t,K-1}\right|z\right),L_{t}\left(\left.c_{t,K}\right|z\right)\right].\end{cases}$ (40)
which again ensures
$F_{t}^{*}\left(\left.c_{t,K-1}\right|z\right)=\overline{L}\left(\left.c_{t,K-1}\right|z\right).$
We now show that the constructions of $F_{t}^{*}\left(\left.c\right|w\right)$ at $c=c_{t,K}$ and $c_{t,K-1}$ in (36) and (40) satisfy the monotonicity requirement of a CDF, i.e.,
$F_{t}^{*}\left(\left.c_{t,K-1}\right|w\right)\leq F_{t}^{*}\left(\left.c_{t,K}\right|w\right).$ (41)
To see this, first notice from (36) and the monotonicity of $F_{t}^{U}$ that
$F_{t}^{*}\left(\left.c_{tK}\right|w\right)\geq F_{t}^{U}\left(\left.c_{t,j_{1}-1}\right|w\right)$ (42)
Next, consider the two cases in (40). If $\overline{L}\left(\left.c_{t,K-1}\right|z\right)\in\left[L_{t}\left(\left.c_{t,K-1}\right|z\right),L_{t}\left(\left.c_{t,K}\right|z\right)\right],$then by the construction and the monotonicity of $F_{t}^{L},F_{t}^{U}$, we have
$F_{t}^{*}\left(\left.c_{t,K-1}\right|w\right)\leq F_{t}^{L}\left(\left.c_{t,K-1}\right|w\right)\leq p_{t}\left(w\right)=F_{t}^{U}\left(\left.c_{t,1}\right|w\right),$
which together with (42) imply (41). If $\overline{L}\left(\left.c_{t,K-1}\right|z\right)\in\left[U_{t}\left(\left.c_{t,j_{2}-1}\right|z\right),U_{t}\left(\left.c_{t,j_{2}}\right|z\right)\right]$, then by the monotonicity of $U_{t}\left(\left.\cdot\right|z\right)$ we know that $j_{2}\leq j_{1}$. If in addition $j_{2}\leq j_{1}-1$, then by the monotonicity of $F_{t}^{U}$ we have
$F_{t}^{*}\left(\left.c_{t,K-1}\right|w\right)\leq F_{t}^{U}\left(\left.c_{t,j_{2}}\right|w\right)\leq F_{t}^{U}\left(\left.c_{t,j_{1}-1}\right|w\right),$
which together with (42) imply (41). Otherwise, we must have $j_{2}=j_{1}$, then by (35), (38) and the monotonicity of $\overline{L}\left(\left.\cdot\right|z\right)$, we must have
$\alpha_{2}\geq\alpha_{1}$
since $\overline{L}\left(\left.c_{t,K-1}\right|z\right)\leq\overline{L}\left(\left.c_{t,K}\right|z\right)$ and $U_{t}\left(\left.c_{t,j_{2}-1}\right|z\right)\leq U_{t}\left(\left.c_{t,j_{2}}\right|z\right)$ Therefore, by (36), (40), and the monotonicity of $F_{t}^{U}$, we have
$F_{t}^{*}\left(\left.c_{t,K-1}\right|w\right)$ $=\alpha_{2}F_{t}^{U}\left(\left.c_{t,j_{2}-1}\right|w\right)+\left(1-\alpha_{2}\right)F_{t}^{U}\left(\left.c_{t,j_{2}}\right|w\right)$
$\leq\alpha_{1}F_{t}^{U}\left(\left.c_{t,j_{1}-1}\right|w\right)+\left(1-\alpha_{1}\right)F_{t}^{U}\left(\left.c_{t,j_{2}}\right|w\right)=F_{t}^{*}\left(\left.c_{t,K}\right|w\right).$
(iii) The construction of $F_{t}^{*}\left(\left.c\right|w\right)$ at $c=c_{t,1},...,c_{t,K-2}$ can be carried out in the same way as $F_{t}^{*}\left(\left.c_{t,K-1}\right|w\right)$, and is thus omitted here. In summary, $F_{t}^{*}$ is constructed so that
$F_{t}^{*}\left(\left.c\right|z\right)=\overline{L}\left(\left.c\right|z\right),\quad\forall c\in{\cal C}_{t}$
Furthermore, since for each $c\in{\cal C}_{t}$, the value $F_{t}^{*}\left(\left.c\right|w\right)$ is always constructed as a weighted average of (two) values in
$\left\{F^{L}\left(\left.c\right|w\right),F^{U}\left(\left.c\right|w\right):\ c\in{\cal C}\right\},$
by (32), we have
$F_{t}^{*}\left(\left.c\right|w\right)=p_{t}\left(w\right),\ \forall c\in{\cal C}_{t}.$
Hence, for any $w$, we must have $w_{t}^{{}^{\prime}}\theta\in{\cal C}_{t}$ and thus
$F_{t}^{*}\left(\left.w_{t}^{{}^{\prime}}\theta\right|w\right)=p_{t}\left(w\right),$
which ensures (30).
Step 3:
We now show how to construct $F_{t}^{*}\left(\left.c\right|w\right)$ at $c\in{\cal C}_{-t}:={\cal C}\backslash{\cal C}_{t}$ to ensure (31), using based on the previously assigned values of $F_{t}^{*}\left(\left.c\right|w\right)$ for $c\in{\cal C}_{t}$ in Step 2. We rank-order elements in ${\cal C}_{-t}$ in strict ascending order as follows
$c_{-t,1}<...<c_{-t,\overline{K}}\text{ for some }\overline{K}\geq 1.$
(i) We start with the largest element $c_{-t,\overline{K}}$. Then $\overline{L}\left(\left.c_{-t,\overline{K}}\right|z\right)$ can be expressed as follows in three different scenarios:
@@
$\overline{L}\left(\left.c_{-t,\overline{K}}\right|z\right):=\begin{cases}\alpha_{3}\overline{L}\left(\left.c_{t,K}\right|z\right)+\left(1-\alpha_{3}\right)\cdot 1,&\text{if }c_{-t,\overline{K}}>c_{t,K},\\ \alpha_{4}\overline{L}\left(\left.c_{t,j-1}\right|z\right)+\left(1-\alpha_{3}\right)\overline{L}\left(\left.c_{t,j}\right|z\right),&\text{if }c_{-t,\overline{K}}\in\left(c_{t,j-1},c_{t,j}\right)\text{ for some }j,\\ \alpha_{5}\cdot 0+\left(1-\alpha_{5}\right)\overline{L}\left(\left.c_{t,1}\right|z\right)&\text{if }c_{-t,\overline{K}}<c_{t,1},\end{cases}$
for some $\alpha_{3},\alpha_{4},\alpha_{5}\in\left[0,1\right]$. Accordingly, we can set